OpenStudy (anonymous):

how would I Evaluate the following: a) 1/ 5^-2 b) (-3)^-3

OpenStudy (lalaly):

1/5^-2 = 5^2 = 25 -3^-3 = 1/(-3)^3 = -1/9

OpenStudy (anonymous):

OHH thx so much!!!